### v2

# DehazeGS: Seeing Through Fog with 3D Gaussian Splatting

Abstract
Current novel view synthesis tasks primarily rely on high-quality and clear images. However, in foggy scenes, scattering and attenuation can significantly degrade the reconstruction and rendering quality. Although NeRF-based dehazing reconstruction algorithms have been developed, their use of deep fully connected neural networks and per-ray sampling strategies leads to high computational costs. Moreover, NeRF’s implicit representation struggles to recover fine details from hazy scenes. In contrast, recent advancements in 3D Gaussian Splatting achieve high-quality 3D scene reconstruction by explicitly modeling point clouds into 3D Gaussians. In this paper, we propose leveraging the explicit Gaussian representation to explain the foggy image formation process through a physically accurate forward rendering process. We introduce DehazeGS, a method capable of decomposing and rendering a fog-free background from participating media using only muti-view foggy images as input. We model the transmission within each Gaussian distribution to simulate the formation of fog. During this process, we jointly learn the atmospheric light and scattering coefficient while optimizing the Gaussian representation of the hazy scene. In the inference stage, we eliminate the effects of scattering and attenuation on the Gaussians and directly project them onto a 2D plane to obtain a clear view. Experiments on both synthetic and real-world foggy datasets demonstrate that DehazeGS achieves state-of-the-art performance in terms of both rendering quality and computational efficiency.
1 Introduction
In recent years, Neural Radiance Fields (NeRF)  Mildenhall et al. (2021) have leveraged deep fully connected neural networks to implicitly represent 3D scenes and achieve impressive rendering quality through differentiable volumetric rendering techniques Levoy (1990); Max (1995). Subsequent works have focused on enhancing NeRF’s performance, particularly in terms of training speed  Chen et al. (2022); Müller et al. (2022) and rendering efficiency Garbin et al. (2021); Yu et al. (2021). A recent work, 3D Gaussian Splatting (3DGS)  Kerbl et al. (2023), abandons NeRF’s implicit rendering approach and instead represents scenes explicitly by converting point clouds into 3D Gaussians (ellipsoids). Thanks to its tile-based multi-threaded parallel rendering mechanism, 3DGS achieves high-quality real-time rendering.
Figure 1: DehazeGS can generate accurate rendering results for scenes with scattering media $(b)$. By learning disentangled representations of the participating media and the clear scene, it simultaneously recovers the clear scene$(c)$ and obtains accurate depth estimations $(d)$.
However, when capturing scenes containing scattering media, the light received by the detector is disrupted by the medium. The detected light mainly consists of two components: first, light reflected from the surfaces of objects that has been attenuated by particles of medium, and second, ambient light (e.g., sunlight) scattered by the particles of medium. This interference often leads to scenes suffering from low contrast and limited visibility. 3DGSs and NeRFs, most methods are designed for clear media, which is disadvantageous for applications such as autonomous driving and robotic operations under foggy scene conditions. Although some dehazing reconstruction methods based on NeRF  Ramazzina et al. (2023); Li et al. (2023) have been proposed, these approaches either suffer from extremely slow training and rendering speeds or are limited to specific indoor scenes. In contrast, traditional dehazing algorithms in low-level computer vision Zhang and Patel (2018); Yang et al. (2022), typically require large-scale paired or unpaired datasets for training. Moreover, whether single-image or multi-image dehazing, these methods fail to meet the requirements of multi-view consistency necessary for 3D reconstruction. Most methods remain confined to 2d image plane Dong et al. (2020); Zheng et al. (2021); Qin et al. (2020), lacking consideration of 3D spatial information. Currently, there are no effective methods, and the ability to perform physically accurate modeling and to separate scattering in participating media is crucial for imaging and scene understanding tasks.
Figure 2: DehazeGS overview, we first obtain the Gaussian distributions in foggy scenes and perform alpha blending on the transmission of each Gaussian distribution to render the transmission map, which is guided and optimized using DCP and BCP priors. We utilize pseudo-depth maps as prior information for depth estimation when optimizing each input image.
To address the above challenges, we propose DehazeGS, the first differentiable physics-based 3DGS dehazing model. We have pioneered a novel representation of 3DGS for foggy scenes. We conceive the idea of modeling Gaussians as scattering particles in the atmosphere. Our method is primarily divided into three components. 1) We establish a mapping relationship between the Gaussian ellipsoid depth $G(z)$ and the transmission Gaussian $G(t)$. By querying the depth of Gaussians as input to a convolutional neural network, we obtain the transmission Gaussian. The scattering coefficients are stored as the weights of the convolutional neural network. 2) After obtaining the transmission Gaussian, we model the foggy Gaussians using the Atmospheric Scattering Model (ASM) McCartney (1976); Narasimhan and Nayar (2002), in conjunction with the estimated atmospheric light coefficients. Once the foggy Gaussians are rasterized, the estimated fog map can be rendered. Using the real fog map as supervision, we jointly optimize the parameters of the Gaussians while training the neural network. 3) As the scene depth increases, light attenuation becomes more pronounced. The accuracy of depth is crucial for estimating the transmission map, and the accuracy of the transmission map, in turn, impacts the dehazing performance. Therefore, to obtain more precise depth estimation, we apply depth regularization by using pseudo-depth maps (as shown in Figure 2) as priors when optimizing each input image. In addition, to emphasize the restoration of distant details, we incorporated a depth-weighted reconstruction loss. Note that for the transmission Gaussian obtained in 1), we directly perform alpha blending to render the transmission map. We use the transmission map generated by the dark channel prior  He et al. (2010) and the bright channel prior Zhang et al. (2021) as supervision to guide the optimization of the rendered transmission map.
Our contributions can be summarized as follows:
1) We propose the first framework for learning clear 3D Gaussian splatting solely from multi-view foggy images. This framework is capable of learning disentangled representations of participating media and clear scenes, achieving fast, high-quality dehazing reconstruction.
2) We propose a novel representation, where atmospheric scattering and clear content are modeled separately. A transmission is established on each Gaussian distribution. Based on the ASM model, the learnable global atmospheric light is combined with the transmission Gaussian and further applied to the latent clear Gaussian, making the Gaussians represent foggy scenes.
3) A novel prior loss committee is introduced, incorporating relevant prior information to effectively enhance the model’s dehazing performance and achieve more accurate depth estimation.
4) Our method achieves real-time dehazing rendering, with most scenes requiring only 3,000 iterations to reach optimal results, and the training time taking approximately 1 minute. Compared to NeRF-based methods, our approach reduces the time overhead by several hundred times.
2 Related Works
2.1 Dehazing Based on Image Processing
In traditional image processing, there are numerous dehazing algorithms that typically leverage the statistical properties of clear images to estimate the transmission map and atmospheric light. These estimates are then used in the inverse process of the ASM to recover the underlying clear image. One of the most classic algorithms is the Dark Channel Prior (DCP) He et al. (2010) dehazing algorithm, which is based on statistical observations from a large number of outdoor haze-free scenes, the majority of local patches contain some pixels with very low intensity in at least one color channel. Subsequently, many other related algorithms have also been developed, such as the Bright Channel Prior Zhang et al. (2021) and the Color Attenuation Prior Zhu et al. (2015). With the advancement of deep learning, neural networks have gradually been applied to image processing, which can be categorized into physics-based approaches and end-to-end methods. In physics-based algorithms, neural networks are primarily used to estimate the atmospheric light and transmission map of hazy images. For example, DCPDN Zhang and Patel (2018) proposed a densely connected pyramid network capable of jointly estimating atmospheric light and transmission maps. PSD Chen et al. (2021) introduced a dehazing network guided by physical priors. For end-to-end dehazing methods, for example, DehazeNet Cai et al. (2016) utilizes a CNN to learn the mapping relationship between the original hazy images and the corresponding medium transmission maps. Dong et al. (2020) proposed a multi-scale enhanced dehazing network with dense feature fusion, which is based on a U-net architecture and directly takes hazy images as input, processing them through multiple network layers to produce clear images.
2.2 Neural Radiance Fields for Dehazing
NeRF Mildenhall et al. (2021) uses a fully connected neural network to represent a scene, where the input is the 5D coordinates of the sampled points, and the output is the volumetric density and view-dependent radiance (color) at those spatial locations. Finally, the color and density of the sampled points are integrated along the view direction to produce colors on the image plane using volumetric rendering techniques. ScatterNeRF Ramazzina et al. (2023) builds upon NeRF by introducing an additional MLP to learn the opacity of the scattering medium and its associated color. A scattering term is added to the volumetric rendering equation, blending clear and blurred sampled points to render foggy images. Its primary limitation lies in the need for high-density sampling along each ray in space, which requires processing through a deep fully connected network to compute the volumetric density and color of sampled points. DehazeNeRF Chen et al. (2023) extends the volumetric rendering equation by simulating the physical reality of atmospheric scattering. It also combines several regularization strategies to achieve 3D shape reconstruction while removing haze. However, this method is primarily designed for indoor scenes. The network in Dehazing-NeRF Li et al. (2023) that estimates the scattering coefficient and atmospheric light requires pre-training, and its framework cannot achieve joint optimization of scattering and clear content. Additionally, it is limited to synthetic datasets, each containing a single object (e.g., lego and chair).
3 Method
3.1 Atmospheric Scattering Model
The Atmospheric Scattering Model (ASM) McCartney (1976); Narasimhan and Nayar (2002), is a physical model that describes the changes in light as it propagates through the atmosphere due to scattering and absorption. Based on the principles of light propagation, it reveals how particles in the atmosphere affect the quality of captured images. Its mathematical formulation is expressed as:
$I{(x)}=J(x)t(x)+A(1-t(x)).$ (1)
Here, $I$ represents the observed hazy image, $J$ denotes the underlying clear image, and $A$ is the global atmospheric light, which is generated by the scattering of ambient light (typically sunlight) by participating media, represents the background light intensity received by imaging devices. It primarily affects the overall brightness of the image. $t(x)$ represents the transmission map, where $x$ denotes the position of a pixel. The transmission map primarily serves to characterize the attenuation of light as it travels through scattering media (fog) in the atmosphere, directly influencing the accuracy of the dehazing results. $t(x)$ can be further expressed as $t(x)=e^{-\beta d(x)}$, where $\beta$ represents the scattering coefficient and $d(x)$ denotes the scene depth at pixel $x$.
3.2 3D Gaussian Splatting
The 3D Gaussian approach does not rely on neural radiance fields. instead, it represents the scene as a series of 3D Gaussian distributions Yu et al. (2024). Based on the initialized sparse point cloud, a set of 3D Gaussians, defined as $G$, is parameterized by its 3D coordinates $\boldsymbol{x}\in\mathbb{R}^{3}$, 3D covariance $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$, opacity $\alpha\in\mathbb{R}$ and color $c\in\mathbb{R}^{3}$. $c$ is represented by spherical harmonics for view-dependent appearance. The distribution of each Gaussian is defined as:
${\mathcal{G}}(\boldsymbol{x})=e^{-\frac{1}{2}{({\boldsymbol{x}-\boldsymbol{\mu}})}^{\boldsymbol{T}}\boldsymbol{\Sigma}^{-1}(\boldsymbol{x}-\boldsymbol{\mu})}.$ (2)
In order to ensure the positive semi-definite property of the covariance matrix during the optimization, it is further expressed as:
$\boldsymbol{\Sigma}=\boldsymbol{R}\boldsymbol{S}\boldsymbol{S}^{\boldsymbol{T}}\boldsymbol{R}^{\boldsymbol{T}},$ (3)
where $\boldsymbol{R}$ and $\boldsymbol{S}$ denote the rotation and scaling matrix.
3D gaussians are projected into the 2D image space from a given camera pose $\boldsymbol{P}_{\boldsymbol{c}}=\{\boldsymbol{R}_{\boldsymbol{c}}\in\mathbb{R}^{3\times 3},\boldsymbol{~{}}\boldsymbol{t}_{\boldsymbol{c}}\in\mathbb{R}^{3}\}$. Given the viewing transformation $\boldsymbol{W}$ and 3D covariance matrix $\boldsymbol{\Sigma}$, the projected 2D covariance matrix $\boldsymbol{\Sigma}^{\prime}$ is computed using, as described in Zwicker et al. (2001):
$\boldsymbol{\Sigma}^{\prime}=\boldsymbol{J}\boldsymbol{W}\boldsymbol{\Sigma}\boldsymbol{W}^{T}\boldsymbol{J}^{T},$ (4)
where $\boldsymbol{J}$ is the Jacobian of the affine approximation of the projective transformation.
Subsequently, the transformed Gaussians are sorted based on their depth and the sorted Gaussians are rasterized to render pixel values using the following volume rendering equation:
$C={\sum\limits_{i\in N}c_{i}}\alpha^{\prime}_{i}{\prod\limits_{j=1}^{i-1}(1-\alpha^{\prime}_{j})},$ (5)
where $c_{i}$ is the learned color and the $\alpha^{{}^{\prime}}_{i}$ is the multiplication result of the learned opacity $\alpha_{i}$ and the 2D Gaussian.
3.3 DehazeGS
Modeling Gaussian distributions in foggy scenes. Fog is a natural phenomenon caused by the scattering effect of aerosol particles in the atmosphere. When photographing in foggy conditions, the light received by the sensor primarily originates from two sources: the light reflected from the object’s surface, which undergoes scattering and attenuation by the particles, along with the environmental light (such as sunlight), which is also scattered by the particles. As light passes through the scattering medium, it experiences varying degrees of scattering and attenuation, with the attenuation increasing as the scene depth increases (i.e., the transmission rate decreases). Since the scene represented by 3DGS is modeled using 3D Gaussians, we propose directly defining the transmission function $t(x)$ on each Gaussian distribution. The physical meaning of this function is the proportion of light intensity that remains unattenuated after passing through the particles. Furthermore, the transmission function is associated with the depth of the Gaussian. For any Gaussian $\mathcal{G}(\boldsymbol{x})$, its depth $d_{\mathcal{G}}(\boldsymbol{x})$ (represents the depth $d$ of the Gaussian ellipsoid at coordinate $x$) is queried, and its corresponding transmission Gaussian $t_{\mathcal{G}}(\boldsymbol{x})$ (represents the transmittance $t$ of the Gaussian ellipsoid at coordinate $x$) is obtained through a single-layer one-dimensional convolutional neural network. The mapping relationship can be expressed as $\left.F_{\beta}:{d_{\mathcal{G}}(\boldsymbol{x})}\rightarrow t_{\mathcal{G}}{(\boldsymbol{x})}\right.$, where $\beta$ represents the learnable weight, that is, the scattering coefficient. The detailed structure of the network is shown in the upper-left part of Fig. 2. Since the transmission range is between 0 and 1, with smaller values indicating more severe attenuation of light, we apply a sigmoid ($\sigma$) activation function at the output of the Gaussian transmission. To ensure the non-negativity of the product of depth and scattering coefficient, we apply a ReLU activation function to the exponential term. Finally, based on equation $t(x)=e^{-\beta d(x)}$ from Section 3.1, we can obtain the transmission function formula based on Gaussians as follows,
$t_{\mathcal{G}}{(\boldsymbol{x})}={\sigma(e}^{-max(0,~{}\beta d_{\mathcal{G}}(\boldsymbol{x}))}),$ (6)
The depth of the Gaussian ellipsoid $d_{\mathcal{G}}(\boldsymbol{x})$ is normalized to the range $[0,1]$, and its gradient is detached to prevent the gradients from flowing through it. This approach is designed to prevent the gradient from affecting the latent representation of clear Gaussians, while allowing for better optimization of the network. For atmospheric light estimation, it is generally assumed that the atmospheric light in foggy scenes is global. The learnable global atmospheric light parameter $A_{\mathcal{G}}$ applies to all Gaussians. Consequently, the Gaussian distribution in foggy scenes is represented as follows,
$\mathcal{G}_{f}{(\boldsymbol{x})}=\mathcal{G}_{c}(\boldsymbol{x})t_{\mathcal{G}}(\boldsymbol{x})+A_{\mathcal{G}}(1-t_{\mathcal{G}}(\boldsymbol{x})),$ (7)
where $\mathcal{G}_{f}{(\boldsymbol{x})}$ and $\mathcal{G}_{c}{(\boldsymbol{x})}$ represent the fogged 3D Gaussians and the latent clear 3D Gaussians, respectively.
This approach of directly establishing the transmission function and atmospheric light on the Gaussians effectively avoids the high sampling and rendering costs associated with NeRF-based method Chen et al. (2023); Ramazzina et al. (2023). Moreover, this approach effectively leverages the reconstruction capability in 3DGS, enhancing the structural consistency of the dehazed scene across multiple viewpoints.
After rendering the fog image, the reconstruction loss between the rendered image and the real fog image is calculated:
$\mathcal{L}_{rec}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L_{\textrm{D-SSIM}}}.$ (8)
Transmission Gaussian Optimization. We directly perform alpha blending on the transmission Gaussians $\mathcal{G}(\boldsymbol{x},t)$ to obtain the rendered transmittance map $\mathcal{\hat{T}}(\boldsymbol{P})$:
$\mathcal{\hat{T}}(\boldsymbol{P})={\sum\limits_{i\in N}t_{\mathcal{G}}(\boldsymbol{x})}\alpha^{\prime}_{i}{\prod\limits_{j=1}^{i-1}(1-\alpha^{\prime}_{j})},$ (9)
where $\boldsymbol{P}$ represents pose. We propose to use the Dark Channel Prior (DCP) He et al. (2010) and further consider Bright Channel Prior (BCP) Sun and Guo (2016) algorithms to guide the optimization of our estimated transmission map. The DCP is one of the most well-known dehazing algorithms, leveraging the statistical properties of clear images to estimate the transmission map and atmospheric light. we follow the prior of traditional methods Chen et al. (2021); Golts et al. (2019) and reformulate the prior as a 3DGS-based energy function:
$\begin{split}\mathcal{L}_{DCP}&={\mathcal{{T}}}^{T}(\boldsymbol{P})L\mathcal{{T}}(\boldsymbol{P})+\lambda{\tilde{\mathcal{T}}(\boldsymbol{P})}^{T}\tilde{\mathcal{T}}(\boldsymbol{P}),\\ \tilde{\mathcal{T}}(\boldsymbol{P})&=(\mathcal{T}(\boldsymbol{P})-\mathcal{\hat{T}}{(\boldsymbol{P})}).\\ \end{split}$ (10)
Here, $\mathcal{T}(\boldsymbol{P})$ represents the transmission map under the corresponding pose $\boldsymbol{P}$, estimated by the DCP algorithm. To mitigate the darkening effect caused by DCP, we further introduce BCP, a brightness enhancement algorithm based on the bright channel prior, which helps to enhance the overall brightness of the rendered dehazed images. The loss function is formulated as follows,
$\left.\mathcal{L}_{BCP}=~{}\middle|\middle|\mathcal{T}(\boldsymbol{P})\left.-\mathcal{\hat{T}}{(\boldsymbol{P})}\middle|\right|_{1}\right.$ (11)
where $\mathcal{T}(\boldsymbol{P})$ is estimated from the BCP.
Depth Supervision Loss. In real-world scenarios, as the depth increases, the influence of scattering and attenuation on light becomes more significant, resulting in reduced visibility of distant details. The accuracy of depth estimation is a key factor in obtaining a more precise transmission map. In 3DGS, depth maps can be rendered by performing alpha blending on the depths of Gaussians.
$\hat{D}={\sum\limits_{i\in N}z_{i}}\alpha^{\prime}_{i}{\prod\limits_{j=1}^{i-1}(1-\alpha^{\prime}_{j})}.$ (12)
To guide the optimization of Gaussian ellipsoid depths, we utilize DepthAnything V2 Yang et al. (2024b) to predict the depth of foggy images, generating a pseudo-depth map to guide the optimization of the 3DGS-rendered depth. We found that, in a very small number of scenarios, introducing the Contrast Limited Adaptive Histogram Equalization (CLAHE) can slightly improve performance. However, in most cases, CLAHE is not required, and directly estimating the depth from multi-view foggy images as supervision is sufficient.
$\left.\mathcal{L}_{d}=~{}\middle|\middle|D_{pseudo}\left.-\hat{D}\middle|\right|_{1}\right.$ (13)
Furthermore, to enhance the recovery of distant details, we incorporate a depth-weighted reconstruction loss $\mathcal{L}_{\hat{D}_{recon}}$:
$\left.\mathcal{L}_{d_{rec}}=~{}\middle|\middle|\hat{D}_{detach}\left.\cdot(\hat{I}-I)\middle|\right|_{1}\right.$ (14)
Finally, our total loss is formulated as follows:
$\mathcal{L}=\mathcal{L}_{rec}+\lambda_{D}\mathcal{L}_{DCP}+\lambda_{B}\mathcal{L}_{BCP}+\lambda_{d}\mathcal{L}_{d}+\lambda_{d_{rec}}\mathcal{L}_{d_{rec}},$ (15)
where $\lambda_{D}$, $\lambda_{B}$, $\lambda_{d}$ and $\lambda_{d_{rec}}$ are trade-off weights.
4 Experiments
4.1 Experimental Setup
Synthetic Foggy Dataset. We selected four representative scenes from the Mip-NeRF Barron et al. (2021) dataset, including two indoor scenes (bonsai and counter) and two outdoor scenes (garden and stump). For each scene, a small subset of images was reserved as the test set, while the remaining images were processed as follows: we first employed a depth estimation network Yang et al. (2024b) to estimate the ground truth depth of the clear images, then randomly generated scattering coefficients and global atmospheric light for each scene. Finally, the images were fogged based on the principles of the atmospheric scattering model.
Real Foggy Dataset. We used three indoor foggy scenes provided by DehzeNeRF Chen et al. (2023), captured with two professional fog machines and one mobile phone. The three scenes are bear, elephant, and lion. Each scene contains 82, 58, and 68 foggy images and 47, 79, and 47 clear images, respectively. We combined the clear and foggy images from the same scene and input them into COLMAP to obtain the sparse point cloud and corresponding poses.
Baselines. In addition to comparing with the original 3DGS Kerbl et al. (2023), we also compared our method with ScatterNeRF Ramazzina et al. (2023) and DehazeNeRF Chen et al. (2023). Furthermore, since SeaSplat Yang et al. (2024a) (a real-time rendering method for underwater scenes) also has some dehazing reconstruction capabilities, we included it in the comparison as well. Since DehazeNeRF has not released its source code, its qualitative and quantitative results are both sourced from its original paper. Note that the views of the real dataset presented in DehazeNeRF were specifically cropped (removing some edge regions of the scene). To ensure a fair comparison, we also cropped the edge regions of the rendered views for the novel perspectives accordingly.
Training Details. All models, with the exception of DehazeNeRF, whose data is sourced from its original paper, all other models were trained on NVIDIA RTX 4090 GPU. For ScatterNeRF and SeaSplat, we followed the default settings provided in their respective papers. Due to the slower training speed of ScatterNeRF, we utilized three GPUs with a batch size of 2048 for training while keeping the total number of iterations consistent with the paper at 250K. SeaSplat was trained on a single GPU with the number of iterations set to 30K, consistent with the settings in its paper. For our model, training was performed on a single GPU. The number of iterations was set to 30K for synthetic datasets and 3K for real-world datasets. $\lambda_{D}$, $\lambda_{B}$, and $\lambda_{d_{rec}}$ were all set to 0.1. $\lambda_{d}$ utilized a continuous learning rate decay function adapted from JaxNeRF, where the initial weight was set to 1 at step 0, and the final weight was reduced to 0.01 at the specified maximum step, with logarithmically interpolated applied between. The learning rate for the convolutional network was set to $1\times 10^{-6}$, with a weight decay factor of 0.1, decaying over the corresponding number of iterations for synthetic and real scenes.
Figure 3: Qualitative comparison of novel view synthesis results on real datasets, the images of DehazeNeRF are taken from its original paper. Our method exhibits finer texture details and is closer to the ground truth (GT) compared to existing methods. We encourage readers to zoom in to further explore the details further. In the supplementary materials, we present additional results of novel view synthesis.
Table 1: Quantitative comparison for novel view synthesis on the real foggy dataset.
Dataset bear elephant lion Avg Time
Method—Metric PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
ScatterNeRF (250K) 9.53 0.302 0.754 9.8 0.339 0.786 8.21 0.326 0.772 $>$ 16 hours
SeaSplat (30K) 12.13 0.586 0.263 11.09 0.539 0.314 12.16 0.53 0.285 $\sim$ 25 mins
3DGS (30K) 13.09 0.597 0.276 13.28 0.605 0.307 12.01 0.507 0.337 $\sim$ 5 mins
DehazeNeRF (50K) 14.92 0.439 0.341 17.87 0.73 0.15 18.01 0.558 0.259 $>$ 4 hours
Ours (3K) 16.79 0.725 0.173 19.19 0.791 0.113 16.53 0.722 0.163 $\sim$ 1.2 mins
4.2 Quantitative and Qualitative Results
Table 1 and Table 2 present the quantitative comparison results on real hazy datasets and synthetic hazy datasets (The data in DehazeNeRF is sourced from its original paper. However, since it does not provide quantitative metrics for the bear and lion scenes, we opted to calculate them based on the rendered images presented in its paper). From the results, it can be observed that our method outperforms existing approaches in both rendering quality and training time. Specifically, for real-world datasets, our method outperforms existing approaches in rendering details, achieving maximum improvements in PSNR and SSIM by 1.32 and 0.324, respectively, compared to the second-best method. Moreover, it achieves optimal performance within only 3K iterations, requiring approximately 1 minute, whereas NeRF-based methods typically take several hours. Additionally, our method achieves a rendering speed that is tens of times faster than that of NeRF-based methods.
Figure 4: Dehazing rendering results of the synthetic foggy dataset. Our method exhibits finer texture details and is closer to the ground truth (GT) compared to existing methods
For the qualitative comparison results, as shown in Fig. 3 and Fig. 4 (we encourage readers to zoom in to further explore the details further), the dehazing results of SeaSplat exhibit noticeable color anomalies, and are missing regions at the edges. Compared to DehazeNeRF, our reconstruction results outperform in both dehazing quality and the details of near and far scenes. Meanwhile, ScatterNeRF fails to reconstruct the scenes on both datasets. Additionally, our model is capable of achieving dehazing reconstruction in both indoor and outdoor complex scenarios.
Table 2: Quantitative comparison for novel view synthesis on the synthetic foggy dataset.
Dataset garden bonsai counter stump Avg Time
Method—Metric PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
ScatterNeRF (250K) 14.15 0.346 0.609 12.57 0.394 0.689 12.46 0.385 0.751 11.82 0.224 0.771 $>$ 20 hours
3DGS (30K) 12.49 0.62 0.255 12.43 0.609 0.161 13.06 0.621 0.193 11.10 0.516 0.298 $\sim$ 10 mins
SeaSplat (30K) 18.42 0.687 0.275 20.4 0.705 0.121 19.16 0.736 0.183 17.78 0.631 0.293 $\sim$ 30 mins
Ours (30K) 18.72 0.742 0.183 22.19 0.726 0.104 19.31 0.704 0.153 19.10 0.637 0.219 $\sim$ 25 mins
Table 3: Ablation study on the synthetic foggy dataset. Our framework (DehazeGS) is employed starting from $\mathcal{L}_{rec}$.
PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
Vanilla 3DGS 12.27 0.591 0.227
$\mathcal{L}_{rec}$ 19.35 0.675 0.189
$\mathcal{L}_{rec}+\mathcal{L}_{DCP}$ 19.66 0.687 0.173
$\mathcal{L}_{DCP}+\mathcal{L}_{BCP}+\mathcal{L}_{rec}$ 19.74 0.697 0.175
$\mathcal{L}_{DCP}+\mathcal{L}_{BCP}+\mathcal{L}_{rec}+\mathcal{L}_{d}$ 19.77 0.692 0.166
All losses 19.83 0.702 0.164
4.3 Ablation Study
In this section, we perform ablation studies on the various components of our framework to evaluate the role and contribution of each part. Specifically, we investigate the effects of depth regularization, depth-weighted, and relevant physical priors on the rendering results. The quantitative results are presented in Table 3. Note that Vanilla 3DGS refers to the version where the modeling of Gaussian distributions for foggy scenes is omitted from our framework (the top left section of Fig. 2), and only the reconstruction loss is used (The results of its novel view synthesis are shown as 3DGS in Fig. 3 and Fig. 4). Starting from $\mathcal{L}_{rec}$, it represents our framework, but with other loss functions removed, retaining only the reconstruction loss. We can see that when all the losses are combined, our total loss function achieves the maximum benefit. However, using only partial loss functions significantly reduces rendering quality.
5 Conclusion
We propose the first framework for learning clear 3D Gaussian splatting solely from multi-view hazy images. Our approach leverages explicit Gaussian representation to interpret the formation of hazy images through a physically accurate forward rendering process. This enables joint optimization of 3D scene representation while learning the participating medium. Overall, our method outperforms existing methods in terms of rendering quality, training speed, and rendering efficiency. However, there is still room for improvement. Although our method can reconstruct finer texture details at a faster speed and achieve better dehazing results, the dehazing of distant scenes still needs further enhancement in certain scenarios. We hope that our approach can inspire future work in related areas.